Hi, I am in ALgebra 1 and having a hard time with this problem.
Solve the system of linear equations by any method.
3x-2y=2
x-2/9y=0
So far this is what I have;
3(3x-2y=2)= 9x-6y=6
9(x-2/9y=0)= 9x-2y=0
I subtracted the two equations 4y/4=6/4, y=3/2
I plugged 3/2 into the first equation and got
3x-2(3/2)=2
3x-3=2
+3 each side
divide 3 each side
x= 5/3
I tried to check the equation and the first one worked but the second equation does not. Any help would be great thank you!

Answer by Alan3354(69443)

Solve the system of linear equations by any method.
3x-2y=2
x-2/9y=0
So far this is what I have;
3(3x-2y=2)= 9x-6y=6
9(x-2/9y=0)= 9x-2y=0
I subtracted the two equations 4y/4=6/4, y=3/2
********************** Should be -4y = 6
y = -3/2
=================================================